OpenStudy (taylor0402):

I need help on 2 questions...I will give a medal

OpenStudy (taylor0402):

What is the best approximation of the solution to the system, to the nearest integer values?

OpenStudy (taylor0402):

i think the answer is (1,-3) @igreen

OpenStudy (anonymous):

yup it is... @Taylor0402

OpenStudy (taylor0402):

can you help me on one more please @MARC_

OpenStudy (anonymous):

i will try my best

OpenStudy (taylor0402):

what is the solution to the system of equations 5x-3y=-9 and 2x-5y=4

OpenStudy (taylor0402):

i am so confused i dont remember what to do

OpenStudy (anonymous):

there are three types of method 1)elimination method 2)substitution method 3)Graphing method

OpenStudy (anonymous):

we will use elimination method bcoz it is faster

OpenStudy (taylor0402):

ok i know i cant use elimination method so that out of the question

OpenStudy (taylor0402):

oops sorry i guess we can than

OpenStudy (anonymous):

\[5x-3y=-9-First~eqn\]\[2x-5y=4-Second~eqn\]

OpenStudy (anonymous):

Multiply the first eqn by 2 and multiply the second eqn by 5

OpenStudy (taylor0402):

ok got it

OpenStudy (anonymous):

in oder to eliminate x

OpenStudy (taylor0402):

so it would be 5(2*5)-3y=-9

OpenStudy (anonymous):

\[10x-6y=-18\]\[10x-25y=20\]

OpenStudy (taylor0402):

ok then what do i do now this is were i always get stuck

OpenStudy (anonymous):

okay,in oder to eliminate x, subtract the first equation with the second eqn

OpenStudy (taylor0402):

????? how would i do that ?????

OpenStudy (anonymous):

\[10x-6y=-18-first~eqn\]\[10x-25y=20-second~eqn\]

OpenStudy (taylor0402):

i got that but would i do this to subtract them

OpenStudy (taylor0402):

\[10x-10x+6y-25\]

OpenStudy (anonymous):